An old pump can drain a pool in 6 hours working alone. A new pump can drain the pool in 4 hours working alone. How long will it take both pumps to empty the pool working together?
Answer by stanbon(75887)

An old pump can drain a pool in 6 hours working alone. A new pump can drain the pool in 4 hours working alone. How long will it take both pumps to empty the pool working together?
-----------------
Old pump rate = 1/6 job/hr
----
New pump rate = 1/4 job/hr
----
Together rate = 1/x job/hr
---------------------------------
Equation:
1/6 + 1/4 = 1/x
2x + 3x = 12
5x = 12
x = 2.4 hrs. = 2 hrs and 24 minutes (time need to drain the pool together)
==========================
Cheers,
Stan H.